Suppose that the functions s and t are defined for all real numbers x as follows.
s(x)=4x+2
t(x)=5x
Write the expressions for (s-t)(x) and (s*t)(x) and evaluate (s+t)(1).

Answer by ikleyn(52796)

.

(s-t)(x) = (4x+2) - 5x = -x + 2.                       ANSWER


(s*t)(x) = (4x+2)*(5x) = 5x*(4x-2) = 20x^2 - 10x.      ANSWER


(s+t)(1) = 4*1 + 2 + 5*1 = 4 + 2 + 5 = 11.             ANSWER

Solved.   //   All questions are answered.